# Inference for Regression with

### 23 Feb 2024 v1

@@ 5.2 CEO Behavior
We re-examine these questions using the Supervised Topic Model with Covariates. To explain CEO behavior, in $\mathbf{g}_{i}$ we include log employment (a measure of firm size) and an indicator for whether the CEO has an MBA degree. To explain sales, in $\mathbf{q}_{i}$ we include log employment and fixed effects for year and country. As before, we use HMC for inference and the same priors for both stratgies. [^20] We impose the anchor word assumption by zeroing out from $\bm{\beta}_{1}$ ($\bm{\beta}_{2}$) the activity that is relatively least likely in Pure Behavior 1 (2). The priors used are the same as in the simulation exercise, except that we set the Dirichlet concentration parameter $\eta=0.1$ to follow the original paper.
As demonstrated both theoretically and through the simulation exercise, the key quantity that governs the relative importance of sampling error and measurement error is $\kappa$. In the context of the CEO behavior data, the empirical analog of $\kappa$ is the product of the square root of the number of observations (CEOs) and the average value of the inverse of the number of activities per CEO. This value is 0.44, which is close to the lowest value of $\kappa$ in the simulation exercise. This suggests that the two-step approach should perform relatively well in this application. To further test our theory, we also estimate the model using data where we first sample 10% of the activities for each CEO, without replacement. This scenario could represent a researcher observing only half of a workday for each CEO, instead of a full five-day workweek. Such sampling increases the analogue of $\kappa$ to 4.26, which is near the highest value of $\kappa$ in the simulation exercise, indicating that we should expect the two-step approach to perform poorly under these conditions.
@@
Suppliers 0.41 0.44 0.32
Production 0.41 0.32 0.46
Just Outsiders 0.72 1.37 0.58
Communication 1.54 1.15 1.49
Multi-Function 1.4 1.17 1.9
Insiders and Outsiders 1.9 1.67 1.9
C-suite 21.57 13.01 33.9
@@
What explains the differences in estimates across stratgies? To answer this question, we plot the estimated CEO indices in Figure 2. Panel (a) plots the estimated CEO indices obtained using the full sample, while Panel (b) plots the estimated CEO indices obtained using the 10% subsample. The blue line represents the local polynomial fit (with confidence intervals). The figure shows that when the full sample is used, both strategies find a large number of CEOs with $\hat{\theta}_{i}$ close to 0 and 1, and a strong correlation between the two estimates. However, the correlation is much weaker for the 10% subsample, suggesting that there is a large scope for mis-measurement of ${\theta}_{i}$. Interestingly, Proposition 2 suggests that the bias in the two-step estimate of $\bm{\phi}_{1}$ can be severe when $\theta_{i}$ has mass near 0 and 1, as appears to be the case in this dataset. This provides an explanation for why the two-step strategy produces smaller estimates of $\bm{\phi}_{1}$ even in the full dataset.
@@ A.1 Proofs for Section 2
Part (i) is standard and ensures the OLS estimator of $\phi_{1}$ without measurement error is well defined with finite asymptotic variance. Part (ii) is made to simplify technical arguments and can be relaxed, e.g., by controlling the rate at which the distribution of $\theta_{i}$ behaves at the boundary of its support. Finally, part (iii) is the same as Assumption 2(v) and is also made to simplify technical derivations and can be relaxed.
@@
where the first inequality is by the union bound, the second is by Assumption 3(ii), the third is by Chernoff’s inequality for Binomial random variables, and the fourth is because $C_{i}\geq c(\log n)^{1+\epsilon}$ for some $c>0$ and $\theta_{i}\geq\delta$ both hold for all $i$ with probability one by Assumptions 3(ii) and 3(iii). Therefore,
@@ A.2 Proofs for Section 3
Suppose that (5) holds and that $\bm{\theta}_{i}$ and $C_{i}$ are independent. Then
$\mathbb{E}\left[\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]=\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\mathbf{B}+\mathbb{E}\left[\frac{1}{C_{i}}\right]\mathbb{E}\left[\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right)\right].$
In addition, if $\mathbb{E}[Y_{i}|\bm{\theta}_{i}]=\bm{\gamma}^{T}\bm{\theta}_{i}$ and $Y_{i}$ and $\mathbf{x}_{i}$ are independent conditional on $(C_{i},\bm{\theta}_{i})$, then
$\mathbb{E}\left[\hat{\mathbf{p}}_{i}Y_{i}\right]=\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{\gamma}.$
@@
$\mathbb{E}\left[\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]$ $=\mathbb{E}\left[\frac{1}{C_{i}^{2}}\mathbb{E}\left[\left.\mathbf{x}_{i}\!\!\!\;\mathbf{x}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]\right]$
$=\mathbb{E}\left[\frac{1}{C_{i}^{2}}\left(\mathbb{E}\left[\left.\mathbf{x}_{i}\right|C_{i},\bm{\theta}_{i}\right]\mathbb{E}\left[\left.\mathbf{x}_{i}\right|C_{i},\bm{\theta}_{i}\right]^{T}+\mathrm{Var}\left[\left.\mathbf{x}_{i}\right|C_{i},\bm{\theta}_{i}\right]\right)\right]$
$=\mathbb{E}\left[\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}+\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right)\right]$
$=\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\mathbf{B}+\mathbb{E}\left[\frac{1}{C_{i}}\right]\mathbb{E}\left[\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right)\right],$
where the second-last line follows from the mean and variance of the multinomial distribution and the final line is by independence of $\bm{\theta}_{i}$ and $C_{i}$. For the second result, using conditional independence of $Y_{i}$ and $\mathbf{x}_{i}$ given $(C_{i},\bm{\theta}_{i})$, we have
$\mathbb{E}\left[\frac{\mathbf{x}_{i}Y_{i}}{C_{i}}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\left.\frac{\mathbf{x}_{i}}{C_{i}}\right|C_{i},\bm{\theta}_{i}\right]\mathbb{E}\left[\left.Y_{i}\right|C_{i},\bm{\theta}_{i}\right]\right]$
$=\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\mathbb{E}\left[\left.Y_{i}\right|\bm{\theta}_{i}\right]\right]$
$=\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{\gamma},$
as required. ∎
@@
Let $\mathbf{B}$ have full rank and let Assumptions 1(ii) and 1(iii) hold. Then
$\left|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]-\mathbb{E}\left[\frac{1}{C_{i}}\right]\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\right|\to_{p}0.$
@@
$\left|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right)\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right|\to_{p}0$
where $(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}$ exists with probability approaching one by Assumption 1(ii) and because $\mathbf{B}$ has full rank. Each element of $\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}$ is bounded between 00 and $1$, so we may deduce by Chebyshev’s inequality that
$\left|\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}\left[\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\right|\to_{p}0.$
Hence by Assumption 1(ii) and Slutsky’s theorem, we have
$\left|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathbb{E}\left[\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}\right|\to_{p}0.$
The result follows by Lemma 1. ∎
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\to_{p}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]+\mathbb{E}\left[\frac{1}{C_{i}}\right]\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right).$ (19)
For the numerator term, again by Assumption 1(iii), we have
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}Y_{i}=(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}Y_{i}\right)+o_{p}(1),$
where by the LLN and Lemma 1, we have
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}Y_{i}\to_{p}\mathbb{E}\left[\hat{\mathbf{p}}_{i}Y_{i}\right]=\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{\gamma}.$
It follows by Assumption 1(ii) and Slutsky’s theorem that
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}Y_{i}\to_{p}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{\gamma}\,.$ (20)
@@
The first result follows from (19) and (20). Note that the matrix on the right-hand side of (19) is bounded below (in Loewner order) by $\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]$ so its inverse is well defined by Assumption 1(i). The second result follows from the approximation $(\mathbf{A}+\bm{\Delta})^{-1}=\mathbf{A}^{-1}-\mathbf{A}^{-1}\bm{\Delta}\mathbf{A}^{-1}+O(\|\bm{\Delta}\|^{2})$ for $\mathbf{A}$ invertible and $\bm{\Delta}$ small. ∎
Proof of Theorem 2.
First consider the denominator term. By Lemma 2 and condition (7), we have $\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\to_{p}\mathbb{E}[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]$. Hence, by Assumption 2(i),

### 24 Mar 2024 v2

@@ 5.2 CEO Behavior
We re-examine these questions using the Supervised Topic Model with Covariates. To explain CEO behavior, in $\mathbf{g}_{i}$ we include log employment (a measure of firm size) and an indicator for whether the CEO has an MBA degree. To explain sales, in $\mathbf{q}_{i}$ we include log employment and fixed effects for year and country. As before, we use HMC for inference and the same priors for both strategies. [^18] We impose the anchor word assumption by zeroing out from $\bm{\beta}_{1}$ ($\bm{\beta}_{2}$) the activity that is relatively least likely in Pure Behavior 1 (2). The priors used are the same as in the simulation exercise, except that we set the Dirichlet concentration parameter $\eta=0.1$ to follow the original paper.
As demonstrated both theoretically and through the simulation exercise, the key quantity that governs the relative importance of sampling error and measurement error is $\kappa$. In the context of the CEO behavior data, the empirical analog of $\kappa$ is the product of the square root of the number of observations (CEOs) and the average value of the inverse of the number of activities per CEO. This value is 0.44, which is close to the lowest value of $\kappa$ in the simulation exercise. This suggests that the two-step approach should perform relatively well in this application. To further test our theory, we also estimate the model using data where we first sample 10% of the activities for each CEO, without replacement, with the minimum of one. This scenario could represent a researcher observing only half of a workday for each CEO, instead of a full five-day workweek. Such sampling increases the analogue of $\kappa$ to 4.26, which is near the middle value of $\kappa$ in the simulation exercise, indicating that we should expect the two-step approach to perform poorly under these conditions.
@@
Suppliers 0.61 0.74 0.32
Production 0.38 0.33 0.46
Just Outsiders 0.74 1.21 0.58
Communication 1.44 1.23 1.49
Multi-Function 1.35 1.12 1.9
Insiders and Outsiders 1.8 1.83 1.9
C-suite 29.78 16.76 33.9
@@
What explains the differences in estimates across strategies? To answer this question, we plot the estimated CEO indices in Figure 2. Panel (a) plots the estimated CEO indices obtained using the full sample, while Panel (b) plots the estimated CEO indices obtained using the 10% subsample. The blue line represents the local polynomial fit (with confidence intervals). The figure shows that when the full sample is used, both strategies find a large number of CEOs with $\hat{\theta}_{i}$ close to 0 and 1, and a strong correlation between the two estimates. However, the correlation is much weaker for the 10% subsample, suggesting that there is a large scope for mis-measurement of ${\theta}_{i}$. Interestingly, Proposition 2 suggests that the bias in the two-step estimate of $\bm{\phi}_{1}$ can be severe when $\theta_{i}$ has mass near 0 and 1, as appears to be the case in this dataset. This provides an explanation for why the two-step strategy produces smaller estimates of $\bm{\phi}_{1}$ even in the full dataset.
@@ A.1 Proofs for Section 2
Part (i) is standard and ensures the OLS estimator of $\phi_{1}$ without measurement error is well defined with finite asymptotic variance. Part (ii) is made to simplify technical arguments and can be relaxed, e.g., by controlling the rate at which the distribution of $\theta_{i}$ behaves at the boundary of its support. Finally, part (iii) is the same as Assumption 2(vi) and is also made to simplify technical derivations and can be relaxed.
@@
where the first inequality is by the union bound, the second is by Assumption 3(ii), the third is by Chernoff’s inequality for Binomial random variables, and the fourth is because $C_{i}\geq c(\log n)^{1+\epsilon}$ for some $c>0$ and $\theta_{i}\geq\delta$ both hold for all $i$ with probability one by Assumption 3(ii)-(iii). Therefore,
@@ A.2 Proofs for Section 3
Suppose that (5) holds. Then
$\mathbb{E}\left[\left.\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]=\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}+\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right),$
and
$\mathbb{E}\left[\left.(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})^{T}\right|C_{i},\bm{\theta}_{i}\right]=\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right).$
@@
$\mathbb{E}\left[\left.\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]$ $=\frac{1}{C_{i}^{2}}\mathbb{E}\left[\left.\mathbf{x}_{i}\!\!\!\;\mathbf{x}_{i}^{T}\right|C_{i},\bm{\theta}_{i}\right]$
$=\frac{1}{C_{i}^{2}}\left(\mathbb{E}\left[\left.\mathbf{x}_{i}\right|C_{i},\bm{\theta}_{i}\right]\mathbb{E}\left[\left.\mathbf{x}_{i}\right|C_{i},\bm{\theta}_{i}\right]^{T}+\mathrm{Var}\left[\left.\mathbf{x}_{i}\right|C_{i},\bm{\theta}_{i}\right]\right)$
$=\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}+\frac{1}{C_{i}}\left(\mathrm{diag}(\mathbf{B}^{T}\bm{\theta}_{i})-\mathbf{B}^{T}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\mathbf{B}\right),$
where the last line follows from the mean and variance of the multinomial distribution. The second result now follows because $\mathbb{E}\left[\left.\hat{\mathbf{p}}_{i}\right|C_{i},\bm{\theta}_{i}\right]=\mathbf{p}_{i}=\mathbf{B}^{T}\bm{\theta}_{i}$. ∎
@@
Let Assumption 1(i)-(iii) hold. Then
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]-\mathbb{E}\left[\frac{1}{C_{i}}\right]\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\right\|\to_{p}0.$
@@
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right)\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\right\|\to_{p}0$
where $(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}$ exists with probability approaching one by Assumption 1(i)-(ii). Each element of $\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}$ is bounded between 00 and $1$, so we may deduce by Chebyshev’s inequality that
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}\left[\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\right\|\to_{p}0.$
Hence, by Assumption 1(ii) and Slutsky’s theorem, we have
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathbb{E}\left[\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}\right]\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}\right\|\to_{p}0.$
The result follows by Lemma 1 and independence of $C_{i}$ and $\bm{\theta}_{i}$. ∎
Lemma 3.
Let Assumption 1(iv) hold. Then
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\to_{p}\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right],\quad\quad\Bigg{(}\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\Bigg{)}^{-1}\to_{p}\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]^{-1}.$
Proof of Lemma 3.
The first result follows by the law of large numbers (LLN). The second result then follows because the rank condition on $\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ implies $\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]$ has full rank. ∎
Lemma 4.
Let Assumption 1 hold. Then
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\to_{p}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right].$
Proof of Lemma 4.
In view of Assumption 1(iii)-(iv), we have
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\mathbf{q}_{i}^{T}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right)\right\|\\ \leq\left(\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|\right)\times\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{q}_{i}\|\to_{p}0.$
Moreover, $(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\to_{p}(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}$ by Assumption 1(i)-(ii). Let $\mathbf{X}_{i}=\mathbf{q}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbb{E}[\mathbf{q}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}]$ and let $D$ denote the dimension of $\mathbf{q}_{i}$. Then
$\mathbb{E}\left[\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}\right\|^{2}_{F}\right]=\frac{1}{n}\sum_{j=1}^{D}\sum_{k=1}^{V}\mathbb{E}\left[\left(\mathbf{X}_{i}\right)_{j,k}^{2}\right]$ $\leq\frac{1}{n}\sum_{j=1}^{D}\sum_{k=1}^{V}\mathbb{E}\left[\left(\mathbf{q}_{i,j}\right)^{2}\left(\hat{\mathbf{p}}_{i,k}\right)^{2}\right]$
$\leq\frac{1}{n}\mathbb{E}\left[\|\mathbf{q}_{i}\|^{2}\right]\to 0,$
by Assumption 1(iv) and the fact that $\hat{\mathbf{p}}_{i}$ is in the simplex. Hence, $\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}\to_{p}0$. The result follows by Slutsky’s theorem, noting that $\mathbb{E}[\mathbf{q}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}]=\mathbb{E}[\mathbf{q}_{i}\!\!\!\;\bm{\theta}_{i}^{T}]\mathbf{B}$. ∎
In what follows, we let $\mathbf{0}$ denote a conformable matrix of zeros.
Lemma 5.
Let Assumption 1 hold. Then
$\Bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}-\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\\ -\left[\begin{array}[]{cc}\mathbb{E}\left[\frac{1}{C_{i}}\right]\bm{S}\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{array}\right]\Bigg{\|}\to_{p}0.$
Proof of Lemma 5.
Note that
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}=\left[\begin{array}[]{cc}\bm{S}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)\bm{S}^{T}&\bm{S}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right)\\ \left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}\right)\bm{S}^{T}&\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}\end{array}\right].$
The result follows by Lemmas 2, 3, and 4. ∎
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\to_{p}\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\\ +\left[\begin{array}[]{cc}\mathbb{E}\left[\frac{1}{C_{i}}\right]\bm{S}\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}[\bm{\theta}_{i}])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\bm{S}^{T}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{array}\right].$ (19)
For the numerator term, first note that
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}Y_{i}=\left[\begin{array}[]{c}\bm{S}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}Y_{i}\right)\\ \frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}Y_{i}\end{array}\right].$ (20)
For the upper block on the right-hand side, we use Assumption 1(iii) to deduce
$\left\|\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}Y_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}Y_{i}\right)\right\|\to_{p}0.$
We have $(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\to_{p}(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}$ by Assumption 1(i)-(ii). Moreover, by the LLN and independence of $\mathbf{x}_{i}$ and $Y_{i}$ conditional on $(C_{i},\bm{\theta}_{i})$, we have
$\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbf{p}}_{i}Y_{i}\to_{p}\mathbb{E}\left[\hat{\mathbf{p}}_{i}Y_{i}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\left.\frac{\mathbf{x}_{i}}{C_{i}}\right|C_{i},\bm{\theta}_{i}\right]\mathbb{E}\left[\left.Y_{i}\right|C_{i},\bm{\theta}_{i}\right]\right]$
$=\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\mathbb{E}\left[\left.Y_{i}\right|\bm{\theta}_{i},\mathbf{q}_{i}\right]\right]$
$=\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{S}^{T}\bm{\gamma}+\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]\bm{\alpha},$
Hence,
@@
$\bm{S}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}Y_{i}\right)\to_{p}\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{S}^{T}\bm{\gamma}+\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]\bm{\alpha}\,.$
Similarly, for the lower block on the right-hand side of (20), we have by the LLN that
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}Y_{i}\to_{p}\mathbb{E}[\mathbf{q}_{i}Y_{i}]=\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\bm{S}^{T}\bm{\gamma}+\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]\bm{\alpha}.$
Combining the above two displays, we have
$\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}Y_{i}\to_{p}\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]\bm{\psi}.$ (21)
@@
The first result follows from (19) and (21). Note that the matrix on the right-hand side of (19) is bounded below (in Loewner order) by $\mathbb{E}\left[\bm{\xi}_{i}\!\!\!\;\bm{\xi}_{i}^{T}\right]$ so its inverse is well defined by Assumption 1(iv). The second result follows because $(\mathbf{A}+\bm{\Delta})^{-1}=\mathbf{A}^{-1}-\mathbf{A}^{-1}\bm{\Delta}\mathbf{A}^{-1}+O(\|\bm{\Delta}\|^{2})$ for $\mathbf{A}$ invertible and $\bm{\Delta}$ small. ∎
The next three lemmas are used in the proof of Theorem 2. They are derived in the sequence-of-populations asymptotic framework.
Lemma 6.
Let Assumption 2(i)-(iii) hold. Then